# FIER: Fine-Grained and Efficient KV Cache Retrieval for Long-context LLM Inference

### v1

@@ Abstract
The Key-Value (KV) cache reading latency increases significantly with context lengths, hindering the efficiency of long-context LLM inference. To address this, previous works propose retaining a small fraction of KV cache based on token importance. For example, KV eviction uses static heuristics to retain tokens, while KV retrieval dynamically selects query-relevant tokens for more adaptive cache management. However, we observe that important tokens are often sparsely distributed across the long context. This sparsity makes existing page-level KV retrieval inaccurate, as each page may include irrelevant tokens and miss critical ones. In this work, we propose Fier, a Fine-Grained and Efficient KV cache Retrieval method. Fier uses 1-bit quantized keys to estimate the importance of each token, resulting in efficient and precise retrieval. Experiments show that Fier matches full KV performance using only 11% of the cache budget across various long-context tasks, reducing decoding latency by 1.2$\times$ to 1.5$\times$.
@@
Dongwei Wang1, Zijie Liu2, Song Wang3, Yuxin Ren1, Jianing Deng4, Jingtong Hu4, Tianlong Chen2, Huanrui Yang1, 1University of Arizona, 2University of North Carolina at Chapel Hill, 3University of Virginia, 4University of Pittsburgh Correspondence: huanruiyang@arizona.edu
@@ 1 Introduction
To address this issue, previous works have proposed to selectively retain only a subset of KV cache entries based on token importance. Among them, one line of work—known as KV eviction (Fig. 1(b))—focuses on retaining fixed-position tokens that typically receive higher attention weights, such as the initial tokens (due to the attention sink phenomenon) and the most recent tokens Xiao et al. (2023); Zhang et al. (2023); Li et al. (2024); Liu et al. (2023). However, these approaches overlook the dynamic nature of KV criticality, i.e., tokens that are evicted earlier may become important in the future. Their inability to recall evicted tokens often results in degraded performance in multi-round QA or long conversation applications. Motivated by this, another line of work—KV retrieval (Fig. 1(c))—has been proposed to dynamically recall tokens that are most relevant to the current query during generation Tang et al. (2024); Chen et al. (2024).
@@
We evaluate Fier across PG19 Rae et al. (2019), LongBench Bai et al. (2023), and the passkey retrieval benchmarks Peng et al. (2023). The results demonstrate the effectiveness of Fier in both generative and retrieval-focused settings. Experiments show that Fier achieves performance comparable to using the full KV cache while requiring only 11% of the cache budget, and consistently outperforms existing KV eviction and retrieval baselines. Additionally, Fier achieves 1.2$\times$ to 1.5$\times$ decoding speedup across different context lengths on a single RTX 4090 GPU. In summary, we make the following contributions in this work:
@@ 2 Related Work
KV Cache Retrieval. KV retrieval methods, including our proposed Fier, dynamically select tokens relevant to the current query. However, existing approaches like Quest Tang et al. (2024) and ArkVale Chen et al. (2024) retrieve at the page level for efficiency, overlooking the sparse distribution of important tokens. In this paper, we propose a fine-grained, token-level retrieval strategy that maintains efficiency while improving accuracy. This design better captures critical information missed by page-level methods.
@@ 3.2 Fine-grained and Efficient KV Retrieval
where $\mathbf{s}_{i}^{K},\mathbf{z}_{i}^{K}\in\mathbb{R}^{1\times d}$ are the per-channel calibrated scaling and bias vectors specific to the $i$-th key vector. Previous KV quantization methods Zhang et al. (2024) aim to optimize these factors through the calibration process to minimize the impact of quantization on the computed attention score. This objective can be formulated as an $\ell_{2}$ loss:
@@ 3.4 Theoretical Analysis of Efficiency
Beyond retrieval accuracy, the efficiency of the selection stage is also critical for practical deployment. We analyze the key cache load ratio incurred during the selection phase and compare that with Quest. For $\mathbf{K}$ stored in float16, we quantize it to 1-bit with group size $g$. Note that in addition to the 1-bit $\mathbf{K}_{Q}$, each group also needs to store a pair of $(s,z)$ in float16. Therefore, the key cache load ratio will be calculated as:
@@
which decreases with a larger group size. Recall that Quest has a load ratio of $2/L$. For fairness, we set $g=32$, which matches the load ratio of $1/8$ with page size $L=16$ as implemented in the Quest baseline.
@@
6:$\mathbf{K}_{Q}=\left\lfloor\frac{\mathbf{K}-\mathbf{z}^{K}}{\mathbf{s}^{K}}\right\rfloor,\mathbf{K}_{Q}\in\{-1,1\}^{l\times d}$ binary
7:$\tilde{\mathbf{K}}=\mathbf{K}_{Q}\odot\mathbf{s}^{K}+\mathbf{z}^{K}$
@@
9:$\tilde{\mathbf{s}}=\mathbf{q}\cdot\tilde{\mathbf{K}}^{\top}$
@@
13:$\mathbf{K}^{\prime}=\mathbf{K}[\mathcal{S}_{q}],\mathbf{V}^{\prime}=\mathbf{V}[\mathcal{S}_{q}]$
@@ 4.1 Setting
Datasets. We evaluate Fier on the language modeling task PG19 Rae et al. (2019). To assess its performance on long-context QA, we further conduct experiments on LongBench Bai et al. (2023) using six representative datasets: NarrativeQA Kočiskỳ et al. (2018), HotpotQA Yang et al. (2018), Qasper Dasigi et al. (2021), TriviaQA Joshi et al. (2017), GovReport Huang et al. (2021), and MultifieldQABai et al. (2023). The detailed information about the six datasets is in Appendix A. We evaluate Fier on the passkey retrieval task Peng et al. (2023) to assess its ability to model long-range dependencies. We also compare responses from Fier and Quest enabled chatbots in Appendix C.
@@ 4.3 Performance Evaluation
4.3.1 PG19 Results
We begin by evaluating language modeling perplexity on PG19, a benchmark consisting of 100 books with an average length of 70k tokens. We evaluate three different models by feeding them texts of varying lengths and compare the results against both Full KV and Quest. Note that both Fier and Quest are evaluated under the same KV cache budget of 4096 tokens. As shown in Fig. 5, Fier achieves performance close to that of Full KV and significantly outperforms Quest on both the LLaMA and Mistral models.
@@ 4.3.3 Passkey Retrieval
We further evaluate Fier’s ability to handle long-distance dependencies. Specifically, we employ the passkey retrieval task Peng et al. (2023) as our benchmark. This task assesses whether the model can retrieve a simple passkey from a large amount of irrelevant content. Following the setup in Tang et al. (2024), we use a context length of 10k tokens for evaluation with both LongChat-7B and a smaller model, LLaMA3-1B. As shown in Tab. 2, KV eviction methods perform poorly due to their inability to recall discarded tokens, while Quest provides noticeable improvements over them. Fier, however, achieves even higher retrieval accuracy, performing well under extremely low budgets of just 32 and 64 tokens, especially improving the long-context processing capability of smaller models.
@@
4.4 Efficiency Profiling
Inference Efficiency. In Fig. 8, we present the decoding latency of 256 tokens on LLaMA-2-7B under different prefill context lengths. To ensure a fair comparison with Full KV, we include both the time spent on computing quantized attention scores and the time required to recall the selected Top-$k$ tokens. We implement the group-wise quantization kernel using Triton, and employ the Top-$k$ CUDA operator to efficiently perform Top-$k$ token recall. Fier’s efficiency gain is mainly attributed to the speedup in the self-attention computation, as it restricts attention to only a small subset of the KV cache. This acceleration becomes more pronounced as the context length increases; for instance, at a context length of 32k tokens, Fier achieves over 1.5× decoding speedup.
@@ 5 Conclusion
We present Fier, a fine-grained and efficient KV cache retrieval algorithm that selects important tokens using 1-bit quantization. By involving all tokens in the computation, Fier enables token-level criticality estimation, leading to improved recall rate and model performance. Extensive experiments across various tasks and model architectures show that Fier consistently outperforms existing methods. Notably, Fier matches full cache performance using only 11% of the KV budget and achieves a 1.2$\times$–1.5$\times$ decoding speedup.
@@ Acknowledgments
This work was based upon High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department. Partial support for this work was provided by NSF grants CNS-2122320, CNS-2133267, and CNS-2328972. Additional support was received through the Amazon Research Award, Cisco Faculty Award, UNC Accelerating AI Awards, NAIRR Pilot Award, OpenAI Researcher Access Award, and the Gemma Academic Program GCP Credit Award.
@@ Appendix B Implementation Details
Appendix C Comparison of Chatbot Responses

### v2

@@ Abstract
[…]
@@
[…]
22footnotetext: Corresponding author
@@ 1 Introduction
To address this issue, previous works have proposed to selectively retain only a subset of KV cache entries based on token importance. Among them, one line of work—known as KV eviction (Fig. 1(b))—focuses on retaining fixed-position tokens that typically receive higher attention weights, such as the initial tokens (due to the attention sink phenomenon) and the most recent tokens Xiao et al. (2023); Zhang et al. (2023); Li et al. (2024); Liu et al. (2023). However, these approaches overlook the dynamic nature of KV criticality, i.e., tokens that are evicted earlier may become important in the future. Their inability to recall evicted tokens often results in degraded performance in multi-round QA or long conversation applications. Motivated by this, another line of work—KV retrieval (Fig. 1(c))—has been proposed to dynamically recall tokens that are most relevant to the current query during generation Tang et al. (2024); Chen et al. (2024a).
@@
[…]
@@ 2 Related Work
KV Cache Retrieval. KV retrieval methods, including our proposed Fier, dynamically select tokens relevant to the current query. However, existing approaches like Quest Tang et al. (2024), ArkVale Chen et al. (2024a), PQCache Zhang et al. (2025), SparQattn Ribar et al. (2023) and MagicPIG Chen et al. (2024b) retrieve at the page or cluster level or compute partial tokens’ criticality for efficiency, overlooking the sparse distribution of important tokens. In this paper, we propose a fine-grained, token-level retrieval strategy that maintains efficiency while improving accuracy. This design better captures critical information missed by existing methods.
@@ 3.2 Fine-grained and Efficient KV Retrieval
where $\mathbf{s}_{i}^{K},\mathbf{z}_{i}^{K}\in\mathbb{R}^{1\times d}$ are the per-channel calibrated scaling and bias vectors specific to the $i$-th key vector.
Previous KV quantization methods Zhang et al. (2024) aim to optimize these factors through the calibration process to minimize the impact of quantization on the computed attention score. This objective can be formulated as an $\ell_{2}$ loss:
@@ 3.4 Theoretical Analysis of Efficiency
Beyond retrieval accuracy, the efficiency of the selection and decoding stage is also critical for practical deployment. We measure the efficiency using the key cache access ratio (CAR), defined as the fraction of the key cache accessed throughout the pipeline. Specifically, we compare both retrieval- and eviction-based methods under the same top-$k$ setting. In the decoding phase, all methods (Fier, Quest, and H2O) access $k$ tokens, so their CAR is identical. The key difference lies in the selection phase. For $\mathbf{K}$ stored in float16, we quantize it to 1-bit with group size $g$. Note that in addition to the 1-bit $\mathbf{K}_{Q}$, each group also needs to store a pair of $(s,z)$ in float16. Therefore, the CAR of Fier will be calculated as:
@@
which decreases with a larger group size. Recall that Quest has a load ratio of $2/L$. For fairness, we set $g=32$, which matches the load ratio of $1/8$ with page size $L=16$ as implemented in the Quest baseline. In contrast, H2O maintains a dynamic pool of size $k/l$ during the selection stage. Therefore, when $k=l/10$, the CAR of the three methods can be summarized in Tab. 1. While retrieval does introduce additional overhead during the selection compared to eviction, its goal is to improve performance with only a modest CAR increase.
@@
[…]
@@
[…]
@@
[…]
@@
Table 1: CAR comparison of different methods (Assuming $k=l/10$).
Method Sel. (%) Dec. (%) Total CAR (%)
H2O 10.0 10.0 20.0
Quest 12.5 10.0 22.5
Fier 12.5 10.0 22.5
@@ 4.1 Setting
Datasets. We evaluate Fier on the language modeling task PG19 Rae et al. (2019). To assess its performance on long-context QA, we further conduct experiments on LongBench Bai et al. (2023) using six representative datasets: NarrativeQA, HotpotQA, Qasper, TriviaQA, GovReport, and MultifieldQA. The detailed information about the six datasets is in Appendix A. We evaluate Fier on the passkey retrieval task Peng et al. (2023) to assess its ability to model long-range dependencies. We also compare responses from Fier and Quest enabled chatbots in Appendix 9.
@@ 4.3 Performance Evaluation
4.3.1 PG19 Results
We begin by evaluating language modeling perplexity on PG19, a benchmark consisting of 100 books with an average length of 70k tokens. We evaluate three different models by feeding them texts of varying lengths and compare the results against both Full KV and Quest (Fig. 5). Note that both Fier and Quest are evaluated under the same KV cache budget of 4096 tokens. Fier achieves performance close to that of Full KV and significantly outperforms Quest on both the LLaMA and Mistral models.
@@ 4.3.3 Passkey Retrieval
We further evaluate Fier’s ability to handle long-distance dependencies. Specifically, we employ the passkey retrieval task Peng et al. (2023) as our benchmark. This task assesses whether the model can retrieve a simple passkey from a large amount of irrelevant content. Following the setup in Tang et al. (2024), we use a context length of 10k tokens for evaluation with both LongChat-7B and a smaller model, LLaMA3-1B. KV eviction methods perform poorly due to their inability to recall discarded tokens, while Quest provides noticeable improvements over them. Fier, however, achieves even higher retrieval accuracy, performing well under extremely low budgets of just 32 and 64 tokens, especially improving the long-context processing capability of smaller models (shown in Tab. 3).
@@ 4.4 Ablation Study
4.5 Efficiency Profiling
Inference Efficiency. In Fig. 8, we present the decoding latency of 256 tokens on LLaMA-2-7B under different prefill context lengths. To ensure a fair comparison with Full KV, we include both the time spent on computing quantized attention and the time required to recall the selected Top-$k$ tokens. We implement the group-wise quantization kernel using Triton, and employ the Top-$k$ CUDA operator to efficiently perform Top-$k$ token recall. Fier’s efficiency gain is mainly attributed to the speedup in the self-attention computation, as it restricts attention to only a small subset of the KV cache. This acceleration becomes more pronounced as the context length increases; for instance, at a context length of 32k tokens, Fier achieves over 1.5× decoding speedup. We also provide a detailed latency breakdown comparison of Fier and Quest in Appendix D.
@@ 5 Conclusion
[…]
@@ Acknowledgments
This work was based upon High Performance Computing (HPC) resources supported by the University of Arizona TRIF, UITS, and Research, Innovation, and Impact (RII) and maintained by the UArizona Research Technologies department and the computational resource supported by TetraMem Inc. Partial support for this work was provided by NSF grants CNS-2122320, CNS-2133267, and CNS-2328972. Additional support was received through the Amazon Research Award, Cisco Faculty Award, UNC Accelerating AI Awards, NAIRR Pilot Award, OpenAI Researcher Access Award, and the Gemma Academic Program GCP Credit Award.
@@ Appendix B Implementation Details
Appendix C Comparison of Fier and MiKV
Different from the methods discussed in the main text, MiKV Yang et al. (2024) adopts a hybrid strategy: during decoding, it stores top tokens in high precision while retaining the remaining tokens in a low-bit format. The main distinction between Fier and MiKV is that MiKV maintains the entire KV cache in mixed precision, whereas Fier only preserves top tokens during decoding. To ensure fairness, we evaluated Fier on the MMLU dataset Hendrycks et al. (2020) using the same cache size as reported for MiKV. Note that the cache size of Fier accounts for the CAR in the selection phase, which is not explicitly stated in the MiKV. As shown in Tab. 6, Fier demonstrates better performance compared to MiKV under the same cache budget.
Table 6: Comparison of Fier and MiKV on MMLU.
Cache Size (%) Method Acc. (%)
25 MiKV 43.9
25 Fier 45.12
20 MiKV 42.7
20 Fier 43.75
Table 7: Time breakdown of Fier and Quest with different Top-$k$.
Stage Fier ($k=2048$) Fier ($k=512$) Quest ($k=2048$)
[…]
End-to-End Latency 20.5ms 16.7ms 18.3ms
GovQA Perf. 34.42 33.98 33.7
Appendix D Latency breakdown comparison with Quest
We provide a detailed latency breakdown between the pipelines of Fier and Quest in Tab. 7. The primary difference lies in the importance estimation stage: Fier uses an extreme 1-bit quantized attention, while Quest adopts page-level attention. In the top-$k$ recall stage: Quest retrieves pages, while Fier directly retrieves top-$k$ tokens. The final top-$k$ self-attention is identical for both methods. In Tab. 7, we compare the average per-layer latency for each token, as well as the end-to-end generation latency. With the same $k$, Fier has higher latency than Quest due to its fine-grained top-$k$ sorting in the recall stage. However, Fier achieves better performance. By reducing $k$ , Fier can achieve lower latency than Quest while still maintaining better performance. Overall, Fier offers a better trade-off between latency and performance compared to Quest.
Appendix E Comparison of Chatbot Responses